OpenStudy (anonymous):

Solve the following system of equations by the substitution method. 10x + 10y = 1 x = y - 3 What is the value of y? -29/20 31/20 -3/2

OpenStudy (anonymous):

Replace x by its value in the first equation to get 10(y-3)+ 10 y =1 Solve this equation for y

OpenStudy (anonymous):

31/20

OpenStudy (anonymous):

Now find x. x= y-3

OpenStudy (anonymous):

10x + 10y = 1 you substitute, ill come out like this 10(y - 3) + 10y = 1 10y - 30 + 10y = 1 10y + 10y = 1 + 30 20y = 31 y = 31/20 All it is is simplifying and arranging it to isolate y, since that is what your looking for.

OpenStudy (anonymous):

-29/20

OpenStudy (anonymous):

x= 31/20 -3=-29/20